RC meeting and Run Plan, Friday Sept. 15, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


  • Hall A call-in instructions:

- Call RC first.
- Techs on site 7-4:30 Mon-Th, 7-11AM Fr. RC will call Hall A work coordinator (Zak) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

  • Hall Access procedure:

- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

Readiness for Beam

  • CW beam is ready
    • MCC completed collimator scan, found edge (https://logbooks.jlab.org/entry/4179721)
    • MCC was able to deliver CW beam with 3He target SOFs with No target
    • Reference cell IC calibration completed with N2 at 22 psi
    • Target polarization is ~33% as of now (experts are monitoring)
    • Shift crews are doing the beam restoration process with Carbon hole
  • New instruction for shift crew when changing target:

Subsystems Checklist

For each subsystem we ask that the responsible person go over the following points:

  1. The overall status of your subsystem
  2. Will your subsystem be ready well before Hall lock-up on Friday or at the latest by Monday morning?
  3. Does your system being ready depend on your team only or are outside resources required?


  • SBS magnet and Bigbite magnets
    • BigBite magnet
    • work as expected
  • Corrector magnets:
    • work as expected
  • Target status (magnets, measurements, laser, installation, alignment, polarization, power etc)
  • HCal (Jiwan Poudel)
    • VME DAC GUI does not read the correct threshold -- pending issue to be fixed
  • Bigbite (Arun Tadepalli)
  • Counting House Alarms (Bill Henry)
  • DAQ (Alex Camsonne)
  • GEMs (Holly Szumila-Vance)
  • Moller polarimeter (Don Jones)
    • Paul K. suggested that we take CW beam (5uA or higher) parity data with IHWP out to check if we have reasonable PITA offsets for that IHWP state.
  • Hodoscopes:

Shift schedule

  • Need TO for Sunday swing

Preliminary run plan

  1. First, bring the beam safely into the hall (Beam recovery procedure: https://sbs.jlab.org/wiki/index.php/HOW_TOs#Beam_Recovery_Procedure)
  2. Target performance assessment by target group, TO
  3. Reference IC calibrations
  1. Take production data if the target is ready
    1. Production: 1h for 5, 10, 20, 30uA => run 30uA until target experts confirm everything is as it should be. These will be singles (not coincidence). PS5 should be set to 2 for helicity readback.
  2. Charge asymmetry: Paul will run feedback once a day until hall C comes up.

List of items in case target isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. Moller measurement
  4. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)


In person: Sanghwa P, Mark J, Jiwan P, Hanjie L, Arun T, Zak. R